The probability that an Avon salesperson sells beauty products to a prospective customer on the first visit to the customer is 0.6. If the salesperson fails to make the seal on the first visit, the probability that the sale will be made on the second visit is 0.3. The salesperson never visits a perspective customer more than twice. What is the probability that the salesperson will make a sale to a particular customer
Answer #1

the probability that the salesperson will not make a sale to a particular customer is = ( 1 - 0,6 )(1 - 0.3) = 0.28

the probability that the salesperson will make a sale to a particular customer is = 1 - 0.28 = 0.72

Answer : 0.72